Here is some information about the



market for onions in Vancouver.



price per unit - 3,5,7,9,11,etc



Quantity demand-81,71,61,51,41etc



Find the equation for the demand curve.



What is the quantity demand when price =8?



What is the value of what is produced when the quantity demand =21?




1
Expert's answer
2022-01-18T17:24:43-0500

Solution:

i.). Equation for the demand curve:

Y = mx + b

Qd = a + bp

Derive b:

b = "\frac{(Q_{2 - }Q_{1} ) }{(P_{2} - P_{1} ) } = \frac{81 - 71}{5 - 3} = \frac{-10}{2} = -5"

b = -5

 

Derive a:

61 = a – 5(7)

61 = a – 35

a = 96

 

Demand equation: Qd = 96 – 5P

 

ii.). Qd = 96 – 5P

Qd when price = 8

Qd = 96 – 5(8) = 96 – 40 = 56

Quantity demanded d when the price is 8 = 56 units

 

iii.). Qd = 96 – 5P

When Qd is = 21

21 = 96 – 5P

5P = 96 – 21

5P = 75

P = 15

The value of what is produced when Qd is 21 = 15